# Cameras as Rays: Pose Estimation via Ray Diffusion

### 22 Feb 2024 v1

@@ 1 Introduction
We have witnessed rapid recent progress toward the goal of obtaining high-fidelity 3D representations given only a sparse set of input images (Zhang et al., 2021; Zhou & Tulsiani, 2023; Long et al., 2022; Cerkezi & Favaro, 2023). However, a crucial requirement is the availability of camera poses corresponding to the 2D input images. This is particularly challenging as structure-from-motion methods fail to reliably infer camera poses under settings with sparsely sampled views (also referred to as sparse-view or wide-baseline in prior works). To fill this performance gap, recent learning-based approaches have examined the task of predicting cameras given a sparse set of input images, and investigated regression (Jiang et al., 2022; Sinha et al., 2023), energy-based modeling (Zhang et al., 2022; Lin et al., 2023) and denoising diffusion (Wang et al., 2023) for inference. However, while exploring a plethora of learning techniques, these methods have largely side-stepped a crucial question: what representation of camera poses should learning-based methods predict?
@@
At first, there may seem to be an obvious answer. After all, every student of projective geometry is taught that (extrinsic) camera matrices are parameterized with a single rotation and a translation. Indeed, all the above mentioned methods adapt this representation (albeit with varying rotation parametrizations e.g., matrices, quaternions, or angles) for predicting camera poses. However, we argue that such a parsimonious global pose representation maybe suboptimal for neural learning, which often benefits from over-parameterized distributed representations. From a geometric perspective, classical bottom-up methods benefit from low-level correspondence across pixels/patches, while learning-based methods that predict global camera representations may not easily benefit from such (implicit or explicit) associations.
@@
In this work, we propose an alternate camera parametrization that recasts the task of pose inference as that of patch-wise ray prediction (Fig. 1). Instead of predicting a global rotation and global translation for each input image, our model predicts a separate ray passing through each patch in each input image. We show that this representation is naturally suited for transformer-based set-to-set inference models that process sets of features extracted from image patches. To recover the camera extrinsics (${\bm{R}}$, ${\bm{t}}$) and intrinsics (${\bm{K}}$) corresponding to a classical perspective camera, we optimize a least-square objective given the predicted ray bundle. It is worth noting that the predicted ray bundle itself can be seen as an encoding of a generic camera as introduced in Grossberg & Nayar (2001); Schops et al. (2020), which can capture non-perspective cameras such as catadioptric imagers whose rays may not even intersect at a center of projection.
@@
We first illustrate the effectiveness of our distributed ray representation by training a patch-based transformer with a standard regression loss. We show that this already surpasses the performance of state-of-the-art pose prediction methods that tend to be far more compute heavy (Lin et al., 2023; Sinha et al., 2023; Wang et al., 2023). However, there are natural ambiguities in the predicted rays due to symmetries and partial observations (Zhang et al., 2022; Wang et al., 2023). We extend our regression-based method to a denoising diffusion-based probabilistic model and find that this further improves the performance and can recover distinct distribution modes. We demonstrate our approach on the CO3D dataset (Reizenstein et al., 2021) where we systematically study performance across seen categories as well as generalization to unseen ones. Moreover, we also show that our approach can generalize even to unseen datasets and present qualitative results on in-the-wild self-captures. In summary, our contributions are as follows:
@@
  • We extend this approach to capture the distribution over cameras by learning a denoising diffusion model over our ray camera parametrization, leading to further performance gains.
@@ 2.1 Structure-from-Motion and SLAM
Both Structure-from-motion and SLAM aim to recover camera poses and scene geometry from a large set of unordered or ordered images. Classic SfM (Snavely et al., 2006) and indirect SLAM (Mur-Artal et al., 2015; Mur-Artal & Tardós, 2017; Campos et al., 2021) methods generally rely on finding correspondences (Lucas & Kanade, 1981) between feature points (Bay et al., 2006; Lowe, 2004) in overlapping images, which are then efficiently optimized (Schönberger & Frahm, 2016; Schönberger et al., 2016) into coherent poses using Bundle Adjustment (Triggs et al., 1999). Subsequent works have improved the quality of features (DeTone et al., 2018), correspondences (Shen et al., 2020; Yang & Ramanan, 2019; Sarlin et al., 2020), and improved the bundle adjustment process itself (Tang & Tan, 2019; Lindenberger et al., 2021). On the contrary, rather than minimize geometric reconstruction errors, indirect SLAM methods (Davison et al., 2007; Schops et al., 2019) generally optimize photometric errors.
While the methods described in this section can achieve (sub)pixel-perfect accuracy, their reliance on dense and large image sets makes them unsuitable for sparse-view pose estimation.
@@ 2.2 Pose Estimation from Sparsely Sampled Views
Most related to our approach are methods that can predict poses given multiple images. RelPose (Zhang et al., 2022) and RelPose++ (Lin et al., 2023) use energy-based models to compose relative rotations into sets of camera poses. SparsePose (Sinha et al., 2023) learns to iteratively refine sparse camera poses given an initial estimate while FORGE (Jiang et al., 2022) exploits synthetic data to learn camera poses. Most comparable to us is PoseDiffusion (Wang et al., 2023), which also uses a diffusion model to denoise camera poses. However, PoseDiffusion denoises the camera parameters directly whereas we denoise camera rays which we demonstrate to be more precise.
@@ 2.3 Ray-based Camera Parameterizations
Some prior work in calibrating generic camera representations has made use of ray-based representations of cameras, mainly for lenses such as fish-eyed for which the pinhole model is not a good approximation (Kannala & Brandt, 2006). Grossberg & Nayar (2001); Dunne et al. (2010) consider the most general camera model, where each pixel projection is modeled by its ray. Even with better calibration algorithms (Schops et al., 2020), the large number of parameters in these camera models makes them difficult to calibrate accurately. While these works also make use of ray-based camera representations, their focus is on calibration (intrinsics) and require known calibration patterns.
@@ 3 Method
We aim to recover cameras from a sparse image set $\{I_{1},\ldots,I_{N}\}$. Rather than predict global camera parametrizations directly as done in previous work, we propose a ray-based representation that can be seamlessly converted to and from the classic representation (Sec. 3.1). We then describe a regression-based architecture for predicting ray-based cameras in Sec. 3.2. We build on this architecture to introduce a probabilistic framework that estimates the rays using diffusion to handle uncertainties and symmetries that arise from sparsely sampled views in Sec. 3.3.
@@ 3.1 Representing Cameras with Rays
Distributed Ray Representation. Typically, a camera is parameterized by its extrinsics (rotation ${\bm{R}}\in\text{SO}(3)$, translation ${\bm{t}}\in{\mathbb{R}}^{3}$) and intrinsics matrix ${\bm{K}}\in{\mathbb{R}}^{3\times 3}$. While this parameterization compactly relates the relationship of world coordinates to pixel coordinates using camera projection (${\bm{u}}={\bm{K}}[{\bm{R}}~{}|~{}{\bm{T}}]{\bm{x}}$), we hypothesize that it may be difficult for a neural network to directly regress this low-dimensional representation. Instead, inspired by generalized camera models (Grossberg & Nayar, 2001; Schops et al., 2020) used for calibration, we propose to over-parameterize a camera as a collection of rays:
@@
where each ray ${\bm{r}}_{i}\in{\mathbb{R}}^{6}$ is associated with a known pixel coordinate ${\bm{u}}_{i}$. We parameterize each ray ${\bm{r}}$ traveling in direction ${\bm{d}}\in{\mathbb{R}}^{3}$ through any point ${\bm{p}}\in{\mathbb{R}}^{3}$ using the Plücker representation (Plücker, 1828):
@@ 3.4 Implementation Details
To predict cameras with our ray denoiser, we use DDPM (Ho et al., 2020) inference with slight modifications. Empirically, we found that removing the stochastic noise in DDPM inference and stopping the backward diffusion process early (and using the predicted $x_{0}$ as the estimate) yielded improved performance. We hypothesize that this is because while the earlier diffusion steps help select among distinct plausible modes, the later steps yield samples around these—and this may be detrimental to accuracy metrics that prefer distribution modes.
@@ 4.1 Experimental Setup
RelPose++ (Lin et al., 2023). RelPose++ builds upon the pairwise scoring network of RelPose to incorporate multi-view reasoning via a transformer and also allows predicting camera translations.
R+T Regression (Lin et al., 2023). To test the importance of modeling uncertainty, Lin et al. (2023) trains a baseline that directly regresses poses. We report the numbers from Lin et al. (2023).

### 19 Mar 2024 v2

@@ 1 Introduction
We have witnessed rapid recent progress toward the goal of obtaining high-fidelity 3D representations given only a sparse set of input images (Zhang et al., 2021; Goel et al., 2022; Zhou & Tulsiani, 2023; Long et al., 2022; Cerkezi & Favaro, 2024; Truong et al., 2023). However, a crucial requirement is the availability of camera poses corresponding to the 2D input images. This is particularly challenging as structure-from-motion methods fail to reliably infer camera poses under settings with sparsely sampled views (also referred to as sparse-view or wide-baseline in prior works). To fill this performance gap, recent learning-based approaches have examined the task of predicting cameras given a sparse set of input images, and investigated regression (Jiang et al., 2024; Sinha et al., 2023), energy-based modeling (Zhang et al., 2022; Lin et al., 2024) and denoising diffusion (Wang et al., 2023) for inference. However, while exploring a plethora of learning techniques, these methods have largely side-stepped a crucial question: what representation of camera poses should learning-based methods predict?
@@
At first, there may seem to be an obvious answer. After all, every student of projective geometry is taught that (extrinsic) camera matrices are parameterized with a single rotation and a translation. Indeed, all of the above-mentioned methods adapt this representation (albeit with varying rotation parametrizations e.g., matrices, quaternions, or angles) for predicting camera poses. However, we argue that such a parsimonious global pose representation maybe suboptimal for neural learning, which often benefits from over-parameterized distributed representations. From a geometric perspective, classical bottom-up methods benefit from low-level correspondence across pixels/patches, while learning-based methods that predict global camera representations may not easily benefit from such (implicit or explicit) associations.
@@
In this work, we propose an alternate camera parametrization that recasts the task of pose inference as that of patch-wise ray prediction (Fig. 1). Instead of predicting a global rotation and global translation for each input image, our model predicts a separate ray passing through each patch in each input image. We show that this representation is naturally suited for transformer-based set-to-set inference models that process sets of features extracted from image patches. To recover the camera extrinsics (${\bm{R}}$, ${\bm{t}}$) and intrinsics (${\bm{K}}$) corresponding to a classical perspective camera, we optimize a least-square objective given the predicted ray bundle. It is worth noting that the predicted ray bundle itself can be seen as an encoding of a generic camera as introduced in Grossberg & Nayar (2001), which can capture non-perspective cameras such as catadioptric imagers or orthographic cameras whose rays may not even intersect at a center of projection.
@@
We first illustrate the effectiveness of our distributed ray representation by training a patch-based transformer with a standard regression loss. We show that this already surpasses the performance of state-of-the-art pose prediction methods that tend to be much more compute-heavy (Lin et al., 2024; Sinha et al., 2023; Wang et al., 2023). However, there are natural ambiguities in the predicted rays due to symmetries and partial observations (Zhang et al., 2022; Wang et al., 2023). We extend our regression-based method to a denoising diffusion-based probabilistic model and find that this further improves the performance and can recover distinct distribution modes. We demonstrate our approach on the CO3D dataset (Reizenstein et al., 2021) where we systematically study performance across seen categories as well as generalization to unseen ones. Moreover, we also show that our approach can generalize even to unseen datasets and present qualitative results on in-the-wild self-captures. In summary, our contributions are as follows:
@@
  • We extend this approach to capture the distribution over cameras by learning a denoising diffusion model over our ray-based camera parametrization, leading to further performance gains.
@@ 2.1 Structure-from-Motion and SLAM
Both Structure-from-motion and SLAM aim to recover camera poses and scene geometry from a large set of unordered or ordered images. Classic SfM (Snavely et al., 2006) and indirect SLAM (Mur-Artal et al., 2015; Mur-Artal & Tardós, 2017; Campos et al., 2021) methods generally rely on finding correspondences (Lucas & Kanade, 1981) between feature points (Bay et al., 2006; Lowe, 2004) in overlapping images, which are then efficiently optimized (Schönberger & Frahm, 2016; Schönberger et al., 2016) into coherent poses using Bundle Adjustment (Triggs et al., 1999). Subsequent work has improved the quality of features (DeTone et al., 2018), correspondences (Shen et al., 2020; Yang & Ramanan, 2019; Sarlin et al., 2020), and the bundle adjustment process itself (Tang & Tan, 2019; Lindenberger et al., 2021). On the contrary, rather than minimize geometric reconstruction errors, indirect SLAM methods (Davison et al., 2007; Schops et al., 2019) generally optimize photometric errors. While the methods described in this section can achieve (sub)pixel-perfect accuracy, their reliance on dense and large image sets makes them unsuitable for sparse-view pose estimation.
@@ 2.2 Pose Estimation from Sparsely Sampled Views
Most related to our approach are methods that can predict poses given multiple images. RelPose (Zhang et al., 2022) and RelPose++ (Lin et al., 2024) use energy-based models to compose relative rotations into sets of camera poses. SparsePose (Sinha et al., 2023) learns to iteratively refine sparse camera poses given an initial estimate, while FORGE (Jiang et al., 2024) exploits synthetic data to learn camera poses. The most comparable to us is PoseDiffusion (Wang et al., 2023), which also uses a diffusion model to denoise camera poses. However, PoseDiffusion denoises the camera parameters directly, whereas we denoise camera rays which we demonstrate to be more precise.
@@ 2.3 Ray-based Camera Parameterizations
Some prior work in calibrating generic camera representations has made use of ray-based representations of cameras, mainly for lenses such as fish-eyed for which the pinhole model is not a good approximation (Kannala & Brandt, 2006). Grossberg & Nayar (2001); Dunne et al. (2010) consider the most general camera model, where each pixel projection is modeled by its ray. Even with better calibration algorithms (Schops et al., 2020), the large number of parameters in these camera models makes them difficult to accurately calibrate. Although these works also make use of ray-based camera representations, their focus is on calibration (intrinsics) and require known calibration patterns. Neural Ray Surfaces (Vasiljevic et al., 2020) also considers learning the poses of generic cameras but does so from video rather than sparse views.
Parameterizing viewpoints using camera rays is also commonly used in the novel view synthesis community. Rather than render a full image at once, the pixel-wise appearance is conditioned per ray (Mildenhall et al., 2020; Sitzmann et al., 2021; Watson et al., 2023) given known cameras. In contrast, we aim to recover the camera itself.
@@ 3 Method
Our aim is to recover cameras from a sparse set of images $\{I_{1},\ldots,I_{N}\}$. Rather than predict global camera parametrizations directly as done in previous work, we propose a ray-based representation that can be seamlessly converted to and from the classic representation (Sec. 3.1). We then describe a regression-based architecture to predict ray-based cameras in Sec. 3.2. We build on this architecture to introduce a probabilistic framework that estimates the rays using diffusion to handle uncertainties and symmetries that arise from sparsely sampled views in Sec. 3.3.
@@ 3.1 Representing Cameras with Rays
Distributed Ray Representation. Typically, a camera is parameterized by its extrinsics (rotation ${\bm{R}}\in\text{SO}(3)$, translation ${\bm{t}}\in{\mathbb{R}}^{3}$) and intrinsics matrix ${\bm{K}}\in{\mathbb{R}}^{3\times 3}$. Although this parameterization compactly relates the relationship of world coordinates to pixel coordinates using camera projection (${\bm{u}}={\bm{K}}[{\bm{R}}~{}|~{}{\bm{T}}]{\bm{x}}$), we hypothesize that it may be difficult for a neural network to directly regress this low-dimensional representation. Instead, inspired by generalized camera models (Grossberg & Nayar, 2001; Schops et al., 2020) used for calibration, we propose to over-parameterize a camera as a collection of rays:
@@
where each ray ${\bm{r}}_{i}\in{\mathbb{R}}^{6}$ is associated with a known pixel coordinate ${\bm{u}}_{i}$. We parameterize each ray ${\bm{r}}$ traveling in direction ${\bm{d}}\in{\mathbb{R}}^{3}$ through any point ${\bm{p}}\in{\mathbb{R}}^{3}$ using Plücker corodinates (Plücker, 1828):
@@ 3.4 Implementation Details
To predict cameras with our ray denoiser, we use DDPM (Ho et al., 2020) inference with slight modifications. Empirically, we found that removing the stochastic noise in DDPM inference and stopping the backward diffusion process early (and using the predicted $x_{0}$ as the estimate) produced better performance. We hypothesize that this is because while the earlier diffusion steps help select among distinct plausible modes, the later steps yield samples around these—and this may be detrimental to accuracy metrics that prefer distribution modes.
@@ 4.1 Experimental Setup
RelPose++ (Lin et al., 2024). RelPose++ builds upon the pairwise scoring network of RelPose to incorporate multi-view reasoning via a transformer and also allows predicting camera translations.
R+T Regression (Lin et al., 2024). To test the importance of modeling uncertainty, Lin et al. (2024) trains a baseline that directly regresses poses. We report the numbers from Lin et al. (2024).